OpenStudy (anonymous):

What is the value of the expression? 2^4 + 3^2 need help 4 and 2 is an exponent

OpenStudy (girlgoyleh.):

25

jimthompson5910 (jim_thompson5910):

2^4 means you multiply the base (2) by itself 4 times (the exponent) 2^4 = 2*2*2*2 = 16

OpenStudy (girlgoyleh.):

then 3^2=3x3

jimthompson5910 (jim_thompson5910):

another example 2^7 = 2*2*2*2*2*2*2 = 128 notice how I have 7 twos being multiplied

OpenStudy (anonymous):

ooooh yess

jimthompson5910 (jim_thompson5910):

so what does 2^5 equal?

OpenStudy (anonymous):

32 ?

jimthompson5910 (jim_thompson5910):

good, how about 3^4 ?

OpenStudy (anonymous):

81

jimthompson5910 (jim_thompson5910):

very good. You seem to understand how exponents work which is great.